OpenStudy (anonymous):

Which of the following is the solution to the equation 8^(y – 2) = 2^y ?

hartnn (hartnn):

is y=3 in options?

OpenStudy (cwrw238):

oh yes - well spotted hartnn

OpenStudy (cwrw238):

8^1 = 2^3

OpenStudy (anonymous):

nice eyeball problem (and solution)

OpenStudy (anonymous):

y = 1 y = 3 y = –3 y = –1

hartnn (hartnn):

y=3 it is :)

OpenStudy (anonymous):

8(y-2)= 2y 8y-16=2y 6y=16 y =8/3 check by plugging into original equations 8(8/3-2) =2(8/3) 64/3 -16 =16/3 64/3 - 48/3 =16/3 16/3 = 16/3 it checks